OpenStudy (anonymous):

Find the wrong step. The problem is 7 * square root of 425 Step 1: 7(squareroot of 400 + square root of 25) Step 2: 7(20+5) Step 3: 7(25) Step 4: 175

OpenStudy (anonymous):

Rewrite the wrong step

OpenStudy (anonymous):

step 1 is wrong

OpenStudy (anonymous):

Step 1: 7( sqrt(400) + sqrt(25) ) = 175

OpenStudy (anonymous):

so what should step 1 be?

OpenStudy (solomonzelman):

H\[\sqrt{a+b}~≠~\sqrt{a}+\sqrt{b}\]